real of the form a + b 2 , a; b ∈ Z form s a ring.

1
Expert's answer
2022-01-10T12:36:17-0500

Solution:

Let "Q[\sqrt{2}]=\{a+b \sqrt{2} \mid a, b \in Q\}" . Note that "Q[\sqrt{2}] \subset R" , and the operations of + and . on "Q[\sqrt{2}]" are the usual + and . of real numbers. Not only is "Q[\sqrt{2}]" closed under + and . but "Q[\sqrt{2}]" is a field (a subfield of R ).


- "Q[\sqrt{2}]" is closed under addition. "(a+b \sqrt{2})+(c+d \sqrt{2})=(a+c)+(b+d) \sqrt{2}" .


- "Q[\sqrt{2}]" is closed under multiplication. "(a+b \sqrt{2})(c+d \sqrt{2})=a c+a d \sqrt{2}+b c \sqrt{2}+ b d \sqrt{2}^{2}=(a c+2 b d)+(a d+b c) \sqrt{2}" .


- Addition is associative and commutative on "Q[\sqrt{2}]", since it is associative and commutative in R.


- Identity for addition: "0=0+0 \sqrt{2} \in Q[\sqrt{2}]."


- Inverses for addition: The inverse of "a+b \sqrt{2}" is "-(a+b \sqrt{2})=-a+-b \sqrt{2} \in Q[\sqrt{2}]" .


- Multiplication is associative and commutative on "Q[\sqrt{2}]" , since it is associative and commutative in R.


- Distributivity holds in "Q[\sqrt{2}]" , since it holds in R.


- Identity for multiplication: "1=1+0 \sqrt{2} \in" "Q[\sqrt{2}]".


- Inverses for multiplication: Given "a, b \in Q" with "a+b \sqrt{2} \neq 0" ,( either "a \neq 0\ or\ b \neq 0" ), we need to find "c, d \in Q" such that "(a+b \sqrt{2})(c+d \sqrt{2})=1" .

Hence, "Q[\sqrt{2}]" is a ring.